OpenStudy (anonymous):

solution to the equation 25^2c =sqrt.5 ^(4c + 16) ? choices are c=-5/2 c=-7/2 c=7/2 c=4

OpenStudy (anonymous):

25^2c = 5^(4c) = 5^(1/2(4c+16)) 4c = 1/2(4c+16) 8c = 4c + 16 4c = 16 c = 4

OpenStudy (anonymous):

thank youu (: